OpenStudy (anonymous):

if the tangent at any point on the curve y=x^3+kx^2+x+5, x belongs to R, makes an acute angle with x-axis, then which of the following is correct?? A)k<1 B)k<3^1/2 C)0

OpenStudy (amistre64):

a poly is one of the most basic function to take a derivative of .....

OpenStudy (anonymous):

but what will i do after finding the derivative???

OpenStudy (amistre64):

an acute angle has a value between 0 and 90, the slope of a line is also the tangent of an anlge

OpenStudy (amistre64):

in other words, slope <=0

OpenStudy (amistre64):

ugh ... slope >= 0

OpenStudy (amistre64):

y= x^3+kx^2+x+5 y'= 3x^2+2kx+1 >= 0

OpenStudy (anonymous):

oh...got it! thnx...

OpenStudy (amistre64):

:) good luck